my using the elimination method solve the following equations by elimination:
x+y=38 175000x+200000y=7200000

Answer by richwmiller(17219)

175000x+200000y=7200000
first divide by 1000 to get rid of those zeroes
175x+200y=7200
x+y=38 multiply by 175
175x+200y=7200
175x+175y=6650 subtract from above equation
25y=650
y=650/25=26
plug in 26 in x+y=38